OpenStudy (anonymous):

(Using the quadratic formula) What are the approximate solutions of 4x2 + 3 = −12x to the nearest hundredth?

Parth (parthkohli):

\( \color{Black}{\Rightarrow 4x^2 + 12x + 3 = 0 }\) is your quadratic equation, okay?

Parth (parthkohli):

\[ \color{Black}{\Rightarrow x = {-b \pm {\sqrt{b^2 - 4ac} } \over 2a} }\]

Parth (parthkohli):

Get it?

OpenStudy (anonymous):

https://www.wolframalpha.com/

Parth (parthkohli):

So, \[x = {-12\pm \sqrt{12^2 - 4(4)(3)} \over 2(4)} \]

mathslover (mathslover):

\[\large{ax^2+bx+c=0}\] \[\large{\alpha ,\beta = \frac{-b\pm \sqrt{b^2-4ac}}{2a}}\] \[\large{\textbf{alpha and beta r the roots of the equation}}\]

Parth (parthkohli):

Nuh! \[\alpha = {-b + \sqrt{b^2 - 4ac} \over 2a}\]\[\beta = {-b - \sqrt{b^2 - 4ac} \over 2a} \]

OpenStudy (anonymous):

okay

mathslover (mathslover):

firstly i will prefer you to rearrange the given equation in the form of standard form of a quadratic equation : \[\large{4x^2+3-12x=0}\] \[\large{4x^2-12x+3=0}\] Compare the second one with \(\large{ax^2+bx+c=0}\) Can you tell me that what is a , b and c ?

mathslover (mathslover):

@simplybeyoutiful ?

mathslover (mathslover):

I will prefer you to have a look at the following tutorials: http://openstudy.com/study#/updates/4ff461cfe4b01c7be8c7b1ea http://openstudy.com/study#/updates/4fe09fa0e4b06e92b86fbb47 http://openstudy.com/study#/updates/4fe08815e4b06e92b86fa074 I am sure that they will help u